OpenStudy (anonymous):

The distance between two cities is one hundred miles, and a woman drives from one city to the other at a rate of fifty mph. At what rate must she return if the total travel time is three hours and forty minutes? 45 mph 54 mph 60 mph 54 mph

OpenStudy (anonymous):

Sorry. I put C.) twice

OpenStudy (anonymous):

60

OpenStudy (anonymous):

How sure are you?

OpenStudy (anonymous):

i am 1% sure

OpenStudy (anonymous):

50*2=100 rt=d r*5/3=100 r=100*3/5 r=60 mph 60*5/3=100 5/3=1 2/3=1 hr 40 min

OpenStudy (anonymous):

simple solution...

OpenStudy (anonymous):

Total Time taken to go from one city to the other is = 100/50 = 2 hours The time left for the return journey is = 1 hour 40 minutes = 5/3 hrs Speed = Distance / time = 100 / (5/3) = 60 mph

OpenStudy (anonymous):

It was an inside joke. Sort of. But thank you

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

Lol its fine. I was just making a joke. Haha

OpenStudy (anonymous):

me too :D